Brendan is twice old as his brother Paolo. Six years from now Brendan will be 1 1/2 times as old as Paolo is then. How old are Paolo and Brendan now?
Answer by TimothyLamb(4379)

x = age brendan now
y = age paolo now
---
x = 2y
x + 6 = 1.5(y + 6)
---
put the system of linear equations into standard form
---
x - 2y = 0
x - 1.5y = 3
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = age brendan now = 12
y = age paolo now = 6
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---